# Constraining the modified friction in gravitational wave propagation with precessing black hole binaries

### v2

@@ III GW generation and luminosity distance measurements
III.1 GWs propagations from BBHs merger and detector network
@@
To date, the LVK collaboration has reported the observation of about 90 confirmed GW events 1 ; 2 . These events arise from the merging of compact binaries including binary black holes, binary neutron stars, and black hole-neutron star binaries. A few of these events may possess spin precession Hannam:2021pit ; Islam:2023zzj . This aspect holds significant potential for advancing research, especially in the realm of GW studies, by refining the accuracy of specific physical parameter determinations Yun:2023ygz . Despite the observational challenges associated with black hole precession, recent methodologies, including the introduction of a precession signal-to-noise ratio 4 have demonstrated the feasibility of detecting BBH precessions. In parallel, various theoretical frameworks now integrate precession into their GW simulations. To simulate the GW signals from precessing BBHs, in this paper, we adopt the phenomenological waveform model IMRPhenomPv3, pioneered by Sebastian Khan et al. This model, IMRPhenomPv3, incorporates a two-spin approach to reflect the latest insights into precession dynamics.
@@
To constrain the modified GW friction using precessing binary black holes with future GW detectors, we focus on the capabilities of a ground-based GW detector network consist of consisting of two LIGO detectors and two third-generation GW detectors (the ET and CE). The specific locations, azimuths of the arms, and lengths of the arms, among other details of the four detectors are summarized in Table 2, see also ref. Muttoni:2023prw for the information of ET and CE. Notably, our configuration utilizes a CE with an arm length of 40 km instead of 20 km, which enables the CE to have a higher detection accuracy. For sensitivity profiles, LIGO is configured with an A+ sensitivity curve, while CE employs the CE2 (Silicon) sensitivity curve 8 , as forecasted by gwinc, and ET utilizes the ET_D curve 7 for assessing strain sensitivity both in amplitude and spectral density. The analysis is grounded on strain data derived from the background noise, calculated using the power spectral density (PSD) for each interferometer. The sampling rate for these interferometers is set at 2048 Hz, which, according to Nyquist’s theorem, establishes the valid signal frequency range as from 0 Hz to half the sampling rate. Additionally, owing to environmental interferences such as ground vibrations and atmospheric pressure fluctuations, which significantly impede detection below 20 Hz, the minimum frequency threshold for our investigation is established at this value.
@@
Cosmic Explorers CE -33.29 149.09 40 km 135.00 45.00 CE2 Silicon
To quantify the influence of precession on the distance estimation of binary black holes, we carry out our research by a specific strategy. As we mentioned, only a few of the GW events in GWTC-3 show possible signals of precession, for example, the events GW200129_065458 Hannam:2021pit , GW190521 LIGOScientific:2020iuh ; LIGOScientific:2020ufj , and GW191109_010717 Zhang:2023fpp . For the third-generation detectors (ET and CE), the detection of precession in binary black holes is expected to be more easy. For a rough and conservatively estimation using the detection rate of precession in GWTC-3 (about 3 in 90), one expects at least 500 precessing events could be detected per year within redshift $z\lesssim 0.5$, considering that about $10^{4}$ GW events can be detected by the third-generation detectors each year for $z\lesssim 0.5$. In this paper, we inject 20 typical precessing binary black hole events for constraining the modified GW friction. These injected datasets are characterized by 15 source parameters $\{m_{1},m_{2},d_{L},\theta_{\rm JN},{\rm ra},{\rm dec},\psi,a_{1},a_{2},\theta_{1},\theta_{2},\phi_{\rm JL},\phi_{12},t_{c},\phi_{c}\}$ where $m_{1}$, $m_{2}$ are the binary black holes’ component masses, $d_{L}$ is the luminosity distance, ${\rm ra}$ and ${\rm dec}$ describe the sky position of the event and $\psi$ is the polarization angle. $\theta_{\rm JN}$ is the inclination angle of the binary system. $a_{1}$, and $a_{2}$ are the dimensionless spin magnitudes of two black holes. The four angles $\theta_{1}$, $\theta_{2}$, $\phi_{\rm JL}$, and $\phi_{12}$ represent the spin misalignment of the binary, which drives the system to precess. $t_{c}$ is the merging time and $\phi_{c}$ is the coalescence phase. Most of the source parameters of these injected signals are the same as the randomly selected 20 GW events in GWTC-3 with redshifts $z\lesssim 0.5$. The main reason for considering this range is that the redshifts of most galaxies in GLADE+ are less than 0.5. The effective precession spin of the precessing binary system is related to the parameters $\theta_{1}$, $\theta_{2}$, $a_{1}$, $a_{1}$, and $q=m_{1}/m_{2}$ as Hannam:2013oca ; Schmidt:2014iyl
@@
In the 20 injected events, we properly choose the injected values of $\theta_{1}$, $\theta_{2}$, $a_{1}$, and $a_{2}$ such that their $\chi_{p}$ is randomly distributed in the range of $[0.3,0.8]$. In our analysis, similar to the treatment in ref. Yun:2023ygz , we marginalize the parameters $t_{c}$ and $\phi_{c}$.
@@ III.2 Estimations of luminosity distances
The aforementioned procedures simulate the sequence of physical phenomena ensuing from the amalgamation of BBHs, encompassing the genesis, transmission, and detection of GWs via a quartet of laser interferometers, each characterized by specific sensitivities. Through the analysis of data harvested from these GW observatories, it becomes feasible to deduce the values of source parameters for each injected event.
We chose to present the results of a typical GW event, the GW190521-like event depicted in Fig. 1, which is believed to exhibit obvious precession LIGOScientific:2020iuh ; LIGOScientific:2020ufj . This injected signal has a precession spin $\chi_{p}=0.7$. For comparison, we additionally inject a similar GW190521-like event but possesses a small precession spin $\chi_{p}=0.1$. Fig. 1 is composed of two corner plots, detailing our parameter estimation results. Specifically, Fig. 1a and  1b elucidate the luminosity distance and orbital inclination parameter estimations derived from the two simulated GW190521-like event with precession spin $\chi_{p}=0.1$ and $\chi_{p}=0.7$, respectively. This event is characterized by the merger of black holes with masses of 98.4 and 57.2 solar masses, respectively, as reported in the latest version of this event in the Gravitational Wave Transient Catalog 2 (GWTC-2) 2 . The orange lines in the plots represent the injected values for comparison, while the values above the box plots provide the estimates for these source parameters. Notably, Fig. 1a (in green) illustrates the scenario with $\chi_{p}=0.1$, and Fig. 1b (in blue) the scenario with $\chi_{p}=0.7$, where the latter exhibits an enhancement in the accuracy of the luminosity distance estimation by approximately 2 times over the less-precessional case. This advancement sets a solid groundwork for subsequent efforts aimed at testing the modified frictions in GW propagations.
Figure 1: Parameter estimation for luminosity distance and the orbital inclination of two injected GW190521-like events: (a) GW190521-like with less precession $\chi_{p}=0.1$; (b) GW190521-like with precession spin $\chi_{p}=0.7$.
@@
IV Redshift distribution of host galaxies
IV.1 The GLADE+ galaxy catalog
@@
The GLADE+ catalog encompasses approximately 22.5 million galaxies and around 750,000 quasars 12 . During data preprocessing, galaxies lacking orientation (right ascension and declination) or redshift information were excluded, resulting in a dataset comprising 21,482,830 entries. However, the representation of high-redshift galaxies in GLADE+ is notably incomplete, rendering its completeness inadequate for the requirements of future third-generation GW detectors 13 . Fig. 2 illustrates the kernel density estimates of the redshifts for all galaxies within the dataset, revealing a significant concentration of GLADE+ galaxies with redshifts spanning from 0.1 to 0.5. To mitigate the impact of catalog data incompleteness on the reliability of the redshift probability distribution for GW sources, we confined our analysis to GW events exhibiting luminosity distances ranging from 900 Mpc to 3000 Mpc. Given the superior precision of redshift data derived from optical observations relative to that of GW information, we assume that the catalog’s orientation and redshift for each galaxy are accurate and devoid of errors.
@@
IV.2 Errorbox construction
From the Bayesian analysis of the injected 20 precessing events using Bilby as described in Sec. III, we obtained both the estimations of the uncertainties of the luminosity distance and sky location for each event. With both the uncertainties, $\Delta\ln(d_{L})$ of luminosity distance and $\Delta\theta^{2}$ of the spatial localization, we define a rectangular error box in the spatial domain where a GW is pinpointed, noting that events at greater luminosity distances correspond to larger error boxes14 ; 17 . Within this framework, we postulate that the redshift of each galaxy within the error box has an equal probability of occurrence 16 .
After locating the error box at the GW source, we meticulously traversed all the data in GLADE+ to identify all conceivable galaxies within the designated error box. The numbers of the host galaxies for the 20 events are within a range of 15 to 2000, a testament to the high-precision spatial localization afforded by the collective prowess of the four-detector array. As shown in ref. Yun:2023ygz the luminosity distance estimates for systems exhibiting precession are significantly more precise compared to their non-precessing counterparts. Consequently, within the precession context, the potential number of host galaxies within the error box can be reduced by a factor of 10, in comparison to non-precessing event Yun:2023ygz . This reduction directly influences the uncertainty associated with the redshift distribution determination for GW events characterized by precession.
V Monte Carlo method (MCMC) for estimation of MODIFIED GW FRICTION

### v3

@@ III GW generation and luminosity distance measurements
III.1 GW propagations from BBH merger and detector network
@@
To date, the LVK Collaboration has reported the observation of about 90 confirmed GW events 1 ; 2 . These events arise from the merging of compact binaries, including binary black holes, binary neutron stars, and black hole-neutron star binaries. A few of these events may possess spin precession Hannam:2021pit ; Islam:2023zzj . This aspect holds significant potential for advancing research, especially in the realm of GW studies, by refining the accuracy of specific physical parameter determinations Yun:2023ygz . Despite the observational challenges associated with black hole precession, recent methodologies, including the introduction of a precession signal-to-noise ratio 4 have demonstrated the feasibility of detecting BBH precessions. In parallel, various theoretical frameworks now integrate precession into their GW simulations. To simulate the GW signals from precessing BBHs, in this paper, we adopt the phenomenological waveform model IMRPhenomPv3 developed in Ref. Khan:2018fmp . This model is built based on its predecessors, such as IMRPhenomC/D, IMRPhenomP, and IMRPhenomPv2 (see Refs. Hannam:2013oca ; Khan:2015jqa ; Husa:2015iqa ; Santamaria:2010yb and references therein) and incorporates a two-spin approach to reflect the latest insights into precession dynamics.
@@
To constrain the modified GW friction using precessing binary black holes with future GW detectors, we focus on the capabilities of a ground-based GW detector network consisting of two LIGO detectors and two third-generation GW detectors (the ET and CE). The specific locations, azimuths of the arms, and lengths of the arms, among other details of the four detectors, are summarized in Table 2, see also Ref. Muttoni:2023prw for information on the ET and CE. Notably, our configuration utilizes a CE with an arm length of 40 km instead of 20 km, which enables the CE to have a higher detection accuracy. For sensitivity profiles, LIGO is configured with an A+ sensitivity curve  [^1] Our purpose for considering two second-generation detectors alongside next-generation detectors (ET and CE) is to use them for improving the accuracy of GW source localization. As ET and CE begin their observations, it is expected that several second-generation detectors will remain operational. Among these, LIGO A+ stands out as it represents an upgrade positioned between Advanced LIGO and third-generation detectors Cooper:2022jfr . Therefore, we included LIGO A+, comprising two detectors, alongside ET and CE in our analysis ., while CE employs the CE2 (Silicon) sensitivity curve 8 , as forecasted by gwinc, and ET utilizes the ET_D curve 7 for assessing strain sensitivity both in amplitude and spectral density. The analysis is grounded on strain data derived from the background noise and calculated using the power spectral density (PSD) for each interferometer. The sampling rate for these interferometers is set at 2048 Hz, which, according to Nyquist’s theorem, establishes the valid signal frequency range as being from 0 Hz to half the sampling rate. Additionally, owing to environmental interferences such as ground vibrations and atmospheric pressure fluctuations, which significantly impede detection below 20 Hz, the minimum frequency threshold for our investigation is established at this value.
Here we would like to mention that our analyses are based on one representative network layout, as mentioned above. We note that while the geometry of the detector network can influence the parameter estimation for individual events, GW sources are expected to be randomly distributed on the sky. As such, when a statistically significant sample of events is considered, the influence of the specific detector configuration is likely to be reduced. The overall combined results should therefore be relatively robust with respect to different network geometries. Exploring the effect of various possible detector configurations in more detail would be an interesting topic for future work.
@@
Cosmic Explorer CE -33.29 149.09 40 km 135.00 45.00 CE2 Silicon
To quantify the influence of precession on the distance estimation of binary black holes, we carry out our research using a specific strategy. As we mentioned, only a few of the GW events in GWTC-3 show possible signals of precession, for example, the events GW200129_065458 Hannam:2021pit , GW190521 LIGOScientific:2020iuh ; LIGOScientific:2020ufj , and GW191109_010717 Zhang:2023fpp . For the third-generation detectors (ET and CE), the detection of precession in binary black holes is expected to be easier. For a rough, conservative estimation using the detection rate of precession in GWTC-3 (about 3 in 90), one expects that at least 500 precessing events could be detected per year within redshift $z\lesssim 0.5$, considering that about $10^{4}$ GW events can be detected by the third-generation detectors each year for $z\lesssim 0.5$. In this paper, we inject 20 typical precessing binary black hole events for constraining the modified GW friction. These injected datasets are characterized by 15 source parameters $\{m_{1},m_{2},d_{L},\theta_{\rm JN},{\rm ra},{\rm dec},\psi,a_{1},a_{2},\theta_{1},\theta_{2},\phi_{\rm JL},\phi_{12},t_{c},\phi_{c}\}$ where $m_{1}$, $m_{2}$ are the binary black holes’ component masses, $d_{L}$ is the luminosity distance, ${\rm ra}$ and ${\rm dec}$ describe the sky position of the event and $\psi$ is the polarization angle. In addition, $\theta_{\rm JN}$ is the inclination angle of the binary system, and $a_{1}$ and $a_{2}$ are the dimensionless spin magnitudes of two black holes. The four angles $\theta_{1}$, $\theta_{2}$, $\phi_{\rm JL}$, and $\phi_{12}$ represent the spin misalignment of the binary, which drives the system to precess. Here $t_{c}$ is the merging time and $\phi_{c}$ is the coalescence phase. Most of the source parameters of these injected signals are the same as the 20 randomly selected GW events in GWTC-3 with redshifts $z\lesssim 0.5$. The main reason for considering this range is that the redshifts of most galaxies in GLADE+ are less than 0.5. The effective precession spin of the precessing binary system is related to the parameters $\theta_{1}$, $\theta_{2}$, $a_{1}$, $a_{1}$, and $q=m_{1}/m_{2}$ as Hannam:2013oca ; Schmidt:2014iyl
@@
In the 20 injected events, we properly choose the injected values of $\theta_{1}$, $\theta_{2}$, $a_{1}$, and $a_{2}$ such that their $\chi_{p}$ is randomly distributed in the range of $[0.3,0.8]$. Additionally, the inclination value $\theta_{\rm JN}$ is randomly sampled within the defined domain to ensure the reliability of the simulation results. In our analysis, similar to the treatment in Ref. Yun:2023ygz , we marginalize the parameters $t_{c}$ and $\phi_{c}$.
@@ III.2 Estimations of luminosity distances
We choose to present the results of a typical GW event, the GW190521-like event depicted in Fig. 1, which is believed to exhibit obvious precession LIGOScientific:2020iuh ; LIGOScientific:2020ufj . For comparison, we additionally injected two events similar to GW190521, which have a small precession spin $\chi_{p}=0.1$ and no precession spin $\chi_{p}=0$, respectively. Figure 1 consists of three angular plots depicting, in detail, the results of our parameter estimation. Specifically, Fig. 1a- 1c elucidate the luminosity distance and orbital inclination parameter estimations derived from the three simulated GW190521-like events with precession spins $\chi_{p}=0$, $\chi_{p}=0.1$, and $\chi_{p}=0.7$, respectively. This event is characterized by the merger of black holes with masses of 98.4 and 57.2 solar masses, respectively, as reported in the latest version of this event in the Gravitational Wave Transient Catalog 2 (GWTC-2) 2 . The orange lines in the plots represent the injected values for comparison, while the values above the box plots provide the estimates for these source parameters. Notably, Fig. 1a (green) illustrates the case with $\chi_{p}=0$, Fig. 1b (green) depicts the case with $\chi_{p}=0.1$, and Fig. 1c (blue) shows the case with $\chi_{p}=0.7$. Among the figures, Fig. 1a shows the lowest accuracy in luminosity distance estimation, with the injected value (represented by the orange line) falling outside the posterior estimate bounds. The accuracy of Fig. 1a is slightly below that of Fig. 1b. In contrast, Fig. 1c achieves roughly twice the accuracy of the less-precessional case shown in Fig. 1b. The introduction of precession contributes to a more accurate estimation of the luminosity distance. This advancement sets a solid groundwork for subsequent efforts aimed at testing the modified frictions in GW propagations.
Figure 1: Parameter estimation for luminosity distance and the orbital inclination of three injected GW190521-like events: (a) GW190521-like with no precession $\chi_{p}=0$; (b) GW190521-like with less precession $\chi_{p}=0.1$; (c) GW190521-like with precession $\chi_{p}=0.7$.
@@
IV Comparison of different detector combinations for parameter estimation
The main purpose of introducing precession in the analysis is to break the degeneracy between $d_{L}$ and $\theta_{\rm JN}$, enabling us to obtain a more precise estimate of the luminosity distance. However, the key challenge lies in measuring precession, which is extremely difficult with second-generation detectors due to their relatively lower sensitivity compared to third-generation detectors. As a result, the uncertainties in the estimates of the 15 source parameters mentioned in Sec. III.2, obtained through Bayesian analysis, are significantly large. Since precession is calculated based on six of these parameters, as described in Eq. (3.1), the uncertainty in precession measurements is consequently also substantial.
Therefore, it is expected that the high sensitivity of third-generation detectors can help one to accurately measure precession. To illustrate this, we simulate a typical case, a GW150921-like signal observed under different detector configurations. Using this case, we investigate which of the three detector combinations, two LIGOs, ET and CE, or all four detectors, can provide the most accurate measurements of luminosity distance and precession. This is a scenario with an injected value of $\chi_{p}=0.7$, and Fig. 2 presents the parameter estimation results for $d_{L}$ and $\theta_{\rm JN}$ in these three configurations.
Figure 2: Parameter estimation for luminosity distance and the orbital inclination of three injected GW190521-like events with precession spin $\chi_{p}=0.7$: (a) GW190521-like detected with 2 LIGO detectors; (b) GW190521-like detected with ET and CE; (c) GW190521-like detected with 2 LIGO, ET and CE.
Fig. 2 consists of three corner plots that provide a detailed overview of our parameter estimation results. Specifically, Fig. 2a, Fig. 2b, and Fig. 2c illustrate the luminosity distance and orbital inclination parameter estimates for three different detector configurations (2 LIGOs, ET and CE, all 4 detectors, respectively) for the same GW event. The orange lines in the plots represent the injected values for comparison, while the values above the box plots provide the estimates for these source parameters.
We can see that the four-detector array configuration yields the best results among the three configurations considered. The 2 LIGO configuration exhibits the largest uncertainty in luminosity distance estimation, approximately three times that of the four-detector array. Although the CE and ET combination shows slightly lower uncertainty, its central estimate deviates significantly from the injected value (3310 Mpc). Consequently, we opted to exclude these two configurations based on a comprehensive evaluation. We also anticipate that the advantages of the four-detector array will become even more pronounced when observing higher redshift targets.
After obtaining the estimation results of 15 source parameters under these three configurations, we calculated the value of $\chi_{p}$ along with its error bars according to Eq. (3.1) as well. Since Eq. (3.1) is not linear and we have a large number of samples of the 6 parameters used to compute $\chi_{p}$, it is appropriate to use Monte Carlo Simulation. The precession spin parameter $\chi_{p}$ was estimated under three different detector configurations for GW190521-like events. For detection with 2 LIGO detectors, we obtained $\chi_{p}=0.74^{+0.20}_{-0.42}$; for detection with ET and CE detectors, $\chi_{p}=0.67^{+0.06}_{-0.04}$; and for detection with 2 LIGO, ET and CE detectors, $\chi_{p}=0.69^{+0.02}_{-0.03}$. It is evident that the measurement of $\chi_{p}$ is significantly more accurate with the four-detector network, achieving a precision that is 12.4 times greater than that of the 2 LIGO configuration and notably higher than that of the ET and CE combination. In the subsequent analysis, we adopt the configuration of the four-detector network.
V Redshift distribution of host galaxies
V.1 The GLADE+ galaxy catalog
@@
The GLADE+ catalog encompasses approximately 22.5 million galaxies and around 750,000 quasars 12 . During data preprocessing, galaxies lacking orientation (right ascension and declination) or redshift information were excluded, resulting in a dataset comprising 21,482,830 entries. However, the representation of high-redshift galaxies in GLADE+ is notably incomplete, rendering its completeness inadequate for the requirements of future third-generation GW detectors 13 ; Yu:2023ico . Fig. 3 illustrates the kernel density estimates of the redshifts for all galaxies within the dataset, revealing a significant concentration of GLADE+ galaxies with redshifts spanning from 0.1 to 0.5. To mitigate the impact of catalog data incompleteness on the reliability of the redshift probability distribution for GW sources, we confined our analysis to GW events exhibiting luminosity distances ranging from 900 Mpc to 3000 Mpc. Although this corresponds to a nominal redshift upper limit of $z\lesssim 0.5$, we emphasize that in our final sample, 17 out of the 20 selected events have redshifts below $z=0.3$, falling well within the redshift range where the GLADE+ catalog maintains reasonable completeness 12 . This selection effectively avoids the regime of severe incompleteness beyond $z\sim 0.3$, as also reflected in the density drop shown in Fig. 3. Given the superior precision of redshift data derived from optical observations relative to that of GW information, we assume that the catalog’s orientation and redshift for each galaxy are accurate and devoid of errors.
@@
V.2 Errorbox construction
From the Bayesian analysis of the injected 20 precessing events using Bilby as described in Sec. III, we obtained both estimations of the uncertainties of the luminosity distance and sky location for each event. With both the uncertainties, $\Delta\ln(d_{L})$ of luminosity distance and $\Delta\theta^{2}$ of the spatial localization, we define a rectangular error box in the spatial domain where a GW is pinpointed, noting that events at greater luminosity distances correspond to larger error boxes.
[…]
We note that our approach implicitly assumes that the posterior distribution in each of these directions is approximately Gaussian, and that the joint distribution factorizes into independent components along the three spatial axes. This approximation enables a computationally efficient statistical cross-matching with galaxy catalogs. Although real posteriors can exhibit skewness or correlations, we find that the cuboid box adequately encloses the high-posterior-density region in most cases. To illustrate our method, we show in Fig. 4 the posterior distribution of Event A in the plane of $\ln d_{L}$ and RA. The red rectangle denotes the $1\sigma$ cuboid error box based on the marginalized standard deviations. We find that although the posterior distribution exhibits mild asymmetry, the box captures the main posterior support region, confirming the practicality of this approach.
Figure 4: Posterior distribution of Event A with $1\sigma$ rectangular error box.
In addition, we provide in Fig. 5 a three-dimensional visualization of the spatial distribution of galaxies within the cuboid error box for Event B. The axes correspond to RA, Dec, and $\ln d_{L}$ (or equivalently redshift), and each black dot denotes a galaxy selected from the GLADE+ catalog that falls within the $1\sigma$ bounds of the posterior uncertainties. This representation serves to emphasize the spatial clustering of galaxies within the localization volume. The elongated structure along the line-of-sight direction is evident and reflects the typically larger uncertainty in luminosity distance compared to sky localization. Such plots confirm that even with approximate box-based modeling, the galaxy distribution within the localization volume retains rich information about the redshift structure, which can be exploited in our statistical inference of modified gravity parameters.
Figure 5: 3D distribution of galaxies in the $1\sigma$ localization volume of a simulated GW event.
Within this framework, we postulate that the redshift of each galaxy within the error box has an equal probability of occurrence 16 . After locating the error box at the GW source, we meticulously traversed all the data in GLADE+ to identify all conceivable galaxies within the designated error box. The numbers of the host galaxies for the 20 events are within a range of 15 to 2000, a testament to the high-precision spatial localization afforded by the collective prowess of the four-detector array.
As shown in Ref. Yun:2023ygz , the luminosity distance estimates for systems exhibiting precession are significantly more precise compared to their non-precessing counterparts. Consequently, within the precession context, the potential number of host galaxies within the error box can be reduced by a factor of 10, in comparison to non-precessing events Yun:2023ygz . This reduction directly influences the uncertainty associated with the redshift distribution determination for GW events characterized by precession.
VI Monte Carlo method (MCMC) for estimation of MODIFIED GW FRICTION